OpenStudy (anonymous):

2. Which equation has a graph perpendicular to the graph of 7x = 14y – 8 ? (1 point) y = –2x – 7 y = –x + 4 y = x + 1 y = 2x + 9

OpenStudy (anonymous):

standard equation of line is y = mx+c put 7x = 14y -8 into standard equation now find m (slope) of the equation Perpendicular line have negative reciprocal slope from the givem line so you will have y = -1/m(x) +c

OpenStudy (anonymous):

thats not an anwser choice

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

I am supposed to help you first. Not to give you answer directly 7x = 14y -8 7x+8 = 14y y = (1/2)x + (4/7) m = 1/2 prependicular slope = -2 y = mx+c y = -2x +c as only one option is there for -2x so its a